OpenStudy (anonymous):

If ON = 9x – 4, LM = 8x + 7, NM = x – 3, and OL = 4y – 8, find the values of x and y for which LMNO must be a parallelogram. A. x = 11, y = 4 B. x = 11, y = 11 C. x = 4, y = 4 D. x = 4, y = 11

Directrix (directrix):

A. x = 11, y = 4 If this is the case, then: 8x + 7 = 9x -4 --> x = 11 4y - 8 = x - 3 --> 4y -8 = 11 -3 and 4y =16 and y = 4.